1. A block is initially moving at 1.80 m/s. It then experiences a uniform acceleration for 1.25 seconds until it reaches a final speed of 8.92 m/s. Assume the block is moving in a straight line, in the same direction for the entire acceleration. Find the magnitude of the acceleration.

2. Two masses move directly toward each other, colliding and bouncing off each other after the collision with no deformation or damage. Which of the following statements is correct?

a. The total momentum of the two combined masses is less than the total momentum of the masses before the collision, while the total KE of the combined masses is the same to the total KE of masses before the collision

b. The total momentum of the two combined masses is the same than the total momentum of the masses before the collision while the total KE of the combined masses is less than the total KE of masses before the collision

c.  The total momentum of the two combined masses is the same than the total momentum of the masses before the collision while the total KE of the combined masses is greater than the total KE of masses before the collision

d.  The total momentum of the two combined masses is less than the total momentum of the masses before the collision while the total KE of the combined masses is less than the total KE of masses before the collision

e.  The total momentum of the two combined masses is the same than the total momentum of the masses before the collision while the total KE of the combined masses is the same than the total KE of masses before the collision
Answer #1

1) Given: initial speed = 1.8 m/s

final speed = 8.92 m/s , time = 1.25s

Using the equation of linear motion we have

v = u + at

8.92 = 1.8 + a*1.25

a = 5.7 m/s^2

so the acceleration is 5.7 m/s^2